# Toward Metaphor-Fluid Conversation Design for Voice User Interfaces

### v1

@@ 4.1.1. Metaphor Identification
Our metaphor collection process drew from two significant sources, combining a comprehensive literature review with user studies to gather a wide range of VUI metaphors. Desai & Twidale (Desai and Twidale, 2023) conducted a literature review following adapted PRISMA guidelines, similar to (Clark et al., 2019). The review focused on peer-reviewed articles from the past decade that discussed VUIs with speech as the primary input and output modality. An initial search in the ACM digital library using a complex query yielded 664 articles, which were screened down to 80 for final analysis. This source also included a user study with 14 participants who interacted with an Amazon Echo Dot and participated in semi-structured interviews. The literature review and user study together identified 93 metaphors related to VUIs. This study is referred as Source A in Figure 2. Similarly, Chin et al. (Chin et al., 2023) contributed through a study (referred to as Source B in Figure 2) in a simulated smart home environment involving 58 participants. Participants interacted with a VUI, completing cognitive tasks (e.g., word puzzles), and were explicitly asked to compare the VUI to other entities using metaphors. This direct elicitation method resulted in 225 metaphors.
Figure 2. A total of 93 metaphors were identified from Source A ((Desai and Twidale, 2023)), and 225 metaphors were collected from Source B ((Chin et al., 2023)). After removing duplicates, 98 unique metaphors remained. From these, 20 metaphors were selected based on their frequency of occurrence and feasibility for practical implementation in VUI design.
@@ 4.1.2. VUI Conversational Use-Contexts
Desai and Twidale (Desai and Twidale, 2023), in a previous study, identified four conversational use-contexts—commands, information seeking, sociality, and error recovery—based on the different types of tasks users perform. Interestingly, although error recovery is not a task in itself, it still leads to conversations and often evokes amusing metaphorical comparisons from users (e.g., a silly child or a pet). Furthermore, they adopt Epley et al.’s (Epley et al., 2007) definition of sociality to include social tasks such as playing games and engaging in small talk—a definition that we also adopt in this work.
While Desai and Twidale (Desai and Twidale, 2023) classified use-contexts through metaphor analysis, other studies have used usage logs to categorize conversational use-contexts. Bentley et al. (Bentley et al., 2018) analyzed usage logs from 88 households over an average of 110 days, identifying ten conversational use-contexts: music, information, automation, small talk, alarms, weather, video, time, lists, and others. Similarly, Kim and Choudhury (Kim and Choudhury, 2021) conducted a 16-week longitudinal study with 12 older adults, identifying eight interaction topics: music, search, basic device control, casual conversation, time, reminders, weather, and others. Ammari et al. (Ammari et al., 2019) analyzed log files from 82 Amazon Alexa and 88 Google Home devices, categorizing interactions into eight themes: search, music, timers, automation, smart home, macros and programming, family interaction, and privacy.
The metaphorical use-contexts identified by Desai and Twidale (Desai and Twidale, 2023) align closely with these findings. The command use-context encompasses several task types reported in prior work, including music, automation, alarms, weather, video, time, and lists (Bentley et al., 2018); music, basic device control, time, and reminders (Kim and Choudhury, 2021); and music, timers, automation, smart home, and macros (Ammari et al., 2019). These tasks share a common interaction pattern: a one-shot call-and-response approach, where users issue a command and expect an immediate response. In some cases, such as automation and device control, a verbal response may not even be necessary.
Similarly, information seeking includes search-related tasks mentioned across studies (Ammari et al., 2019; Bentley et al., 2018; Kim and Choudhury, 2021), while sociality encompasses small talk and casual conversation. The classification is further expanded by introducing error recovery as an additional conversational use-context. Although errors are not an intended feature, they are a persistent part of interactions with commercial VUIs, cutting across all use-contexts. In VUI literature, error recovery is often conceptualized as following unique conversational rules, similar to a difficult conversation (Porcheron et al., 2018; Fischer et al., 2019). Given the high error rate of commercially available VUIs—Kim and Choudhury (Kim and Choudhury, 2021) reported a 33% error rate in their longitudinal study—responding to errors often involves a back-and-forth interaction to clarify misinterpretations and adjust input, making error recovery a distinct aspect of VUI conversations.
@@ 4.2. Measures
To determine the most suitable metaphor aligned with user preferences, we developed a methodology that maps metaphors on a 7-point Likert scale across dimensions of hierarchy and formality. We then assessed the desired levels of hierarchy and professionalism for VUIs in various conversational use-contexts. It is crucial to emphasize that our study’s objective is to identify metaphors that best match users’ mental models for specific VUI tasks. While the metaphors we identify may have broad applicability, designers should ultimately select metaphors that align with their specific design goals.
@@ 4.4. Participants
Our study initially recruited 160 participants through Prolific, [^3] https://www.prolific.com/, a research recruitment platform. However, 30 participants were excluded due to failing attention check questions or timing out, leaving a final sample of 130 participants with an average age of 38.19 years (SD=12.79 years). Selection criteria included a 98% approval rate and at least one year of platform activity. We sought individuals experienced with voice interfaces to ensure meaningful engagement with the study’s tasks and scenarios. To control for cultural variations in metaphor interpretation (Moser, 2000), we limited participation to English-proficient individuals based in the U.S. The participant pool comprised 50.8% females, 46.2% males, and 3.1% non-binary individuals, with 70% holding at least an Associate’s degree. While all participants were U.S. residents and fluent in English, 97% were native speakers. Regarding VUI familiarity, all participants reported at least moderate familiarity, with 75% claiming high or very high familiarity. Usage frequency varied, with all participants using VUIs weekly and 23.1% engaging daily. The survey took an average of 8.5 minutes to complete, and participants received compensation at a fair rate, as informed by Prolific’s policies.
@@ 4.5.1. Mapping Metaphors
Our analysis includes data from all five task scenarios to determine the average formality and hierarchy preferences for each context. The results revealed distinct preferences across different interaction types. When seeking information, participants showed a preference for Voice User Interfaces (VUIs) to exhibit the highest levels of formality (M=4.67) and dominance (M=3.68). In contrast, during social interactions, participants favored a less formal approach from VUIs (M=2.64). Notably, in error recovery situations, participants expressed a preference for VUIs to display the least dominance (M=2.90) while maintaining a slightly formal demeanor (M=4.49). For command-based interactions, participants desired VUIs to maintain neutral formality (M=3.88) and adopt a slightly subordinate posture (M=3.06). Considering all four use-contexts collectively, the ideal VUI, according to participants, should demonstrate neutral formality (M=3.92) and a slightly subordinate stance (M=3.23). These findings suggest a nuanced preference for VUI behavior that adapts to the specific nature of the interaction. Table 2 summarizes means and standard deviations for formality and hierarchy ratings.
@@
Upon further examination of the data, we noticed intriguing patterns in how the metaphors appeared to be grouped. To explore these patterns in greater depth, we performed a cluster analysis. This analysis was not originally part of our research questions, but it provided additional insights into the relationships between metaphors and how they naturally cluster to form meaningful roles.
4.5.2. Clustering Metaphors
To determine the optimal number of clusters, we applied the K-means algorithm to the dataset. The elbow method, depicted in Figure 4(a), suggested a noticeable inflection point at $k=4$, indicating a natural division in the metaphorical space. However, the gradual decline in the within-cluster sum of squares (WCSS) beyond this point suggested that additional meaningful distinctions could be found.
The silhouette scores, shown in Figure 4(b), were used to further evaluate the optimal number of clusters. These scores helped assess the cohesion within each cluster by comparing the similarity of metaphors within the same cluster against others. The scores peaked at $k=4$, closely followed by $k=5$, which indicated that either number could provide valuable groupings.
(a) Elbow method for determining the optimal number of clusters.
(b) Silhouette scores for different numbers of clusters.
Figure 4. Comparison of the Elbow method and Silhouette scores for clustering analysis.
Considering the elbow method and silhouette analysis, we opted to proceed with $k=5$, representing five distinct metaphorical roles. The rationale for this choice was based on examining cluster centroids and their proximity to the use-context means, which showed that one of the centroids in the $k=5$ solution closely aligned with the overall mean across all use-contexts (D=0.29). This indicated that the cluster captured a role that resonated with general user preferences across different scenarios. Moreover, the $k=5$ configuration revealed a clearer separation of roles, including a specific cluster representing neutrality in formality and slight subordination.
Figure 5. K-means clustering of VUI metaphors into five distinct roles based on formality and hierarchy, with centroids indicating the archetypal interaction style for each cluster.
The five clusters represent distinct VUI roles, reflecting different combinations of formality and hierarchy:
  • Companions (Blue Cluster): These VUIs resemble a friendly presence, offering neutral hierarchy and informal interaction. They are well-suited for contexts that benefit from a relaxed and familiar experience, like interactions with a friend. They were closest to Sociality (D=0.80) and furthest to Error Recovery (D=2.43).
  • Facilitators (Yellow Cluster): Facilitators are formal and somewhat dominant, ideal for learning tasks or providing specialized information, akin to roles like a teacher or nurse. Metaphors contained in these roles were closest to Information Seeking (D=1.82) and furthest to Sociality (D=3.46).
  • Entertainers (Green Cluster): These VUIs take on playful, creative roles that may help bring humor and fun into interactions, similar to a pet or child, and are useful for informal, tension-relieving scenarios. Entertainers were adjudged to be the closest to Sociality (D=1.69) and furthest to Information Seeking (D=3.39).
  • Aides (Red Cluster): Aides are formal and slightly subordinate, focused on assisting users in a structured manner, comparable to assistants or butlers. Aides were mapped to be the most appropriate for Information Seeking (D=0.97) and were least appropriate for Sociality (D=2.74).
  • Guides (Purple Cluster): Guides provide expertise and assistance in a neutral, slightly subordinate role, suitable for a variety of tasks, similar to a Genie or search engine. Guides were rated to be the most appropriate for the use-contexts of Commands (D=0.18) and Error Recovery (D=0.80) out of all clusters. They were also relatively close to Sociality (D=1.07) and Information Seeking (D=1.16).
Designing VUIs that align with specific user roles can significantly enhance the user experience by meeting their expectations in different contexts. Figure 5 visualizes these five roles, offering designers guidance for selecting appropriate metaphors based on formality and hierarchy depending on their use-case.
@@ 4.6. Discussion
4.6.1. Non-human and fictional metaphors display user-desired characteristics
We analyzed 20 different metaphors, including both human and non-human characters such as Admirer, Assistant, Boss, Butler, Child, Coach, Companion, Customer Service Agent, Aunt, Flatmate, Friend, Librarian, Nurse, Teacher, and Therapist. Among these, 5 metaphors were distinctly non-human—such as ”Computer” from Star Trek, Encyclopedia, Family Pet, Genie, and Search Engine—and 2 metaphors, specifically ”Computer” from Star Trek and Genie, were fictional in nature. Out of all use-contexts, interestingly, only one human metaphor—Admirer for Sociality—was rated as the most suitable for the chosen use-contexts.
This finding is insightful because it challenges the common assumption that human metaphors would naturally be the most relatable or suitable for VUIs. Instead, our results suggest that non-human and fictional metaphors are often more compatible with users’ desired levels of formality and hierarchy in interactions. This aligns with previous work, such as Pradhan et al. (Pradhan et al., 2019), who found that older adults categorize VUIs as both persons and objects, underscoring the unique hybrid role VUIs occupy. Similarly, Jung et al. (Jung et al., 2022) found no significant preference for human over non-human metaphors in chatbots, challenging the assumption that human-like qualities are inherently more effective. Desai et al. (Desai et al., 2024a) further showed that metaphor preferences are context-dependent, with human metaphors preferred in health but no clear preference in finance.
These findings provide important implications for VUI designers: selecting metaphors that users find appropriate may require moving beyond the human metaphor and instead embracing characters that exist outside of human conventions and cultural associations. There is a natural tendency in conversation design to rely on human metaphors (Desai and Twidale, 2023; Doyle et al., 2019), though these often come with ethical and practical concerns, as previously discussed. By contrast, non-human metaphors remain relatively underutilized, even though previous findings suggest that these metaphors may be better suited for user-centered interaction design (Desai et al., 2024a).
@@ 4.6.1. Non-human and fictional metaphors display user-desired characteristics
The broader role of a Guide, which embodies these characteristics, appears to include a wide range of metaphors—from the human-like Admirer to the more technological Search Engine and the fictional Genie. This flexibility points to a unique user perception where both human and non-human guides are seen as effective, with perhaps a stronger inclination toward non-human entities for roles involving assistance or direction. This observation opens exciting possibilities for future research. By exploring the Guide role, we may be able to uncover additional metaphors that capture similar desirable attributes, potentially creating characters that blend the best elements of both humans and technology.
4.6.2. Exploring the roles of Entertainers and Facilitators
Although we did not identify a specific use-context where metaphors categorized under the Entertainer and Facilitator roles were most suitable, these roles still warrant exploration due to their unique characteristics. Entertainers are marked by low formality and low dominance, whereas Facilitators exhibit high formality and high dominance. In this sense, these roles are diametrically opposed. However, both can have value in specialized contexts, and understanding this requires examining how these metaphors were originally conceptualized. To achieve this, we utilize the metaphor contextualization framework.
The Entertainer role includes metaphors such as ”child” and ”family pet,” which were often applied by users to make sense of conversational challenges or errors. While designing VUIs around the persona of Entertainers may not appeal to everyone, this role can still be useful in alleviating tension during conversational breakdowns. There may be users who prefer a VUI to maintain a casual, subordinate tone regardless of the context. Although Entertainers might not be ideal for all interactions, they present interesting opportunities for VUIs designed to embody lightheartedness and ease.
On the other hand, Facilitators encompass metaphors like ”Boss,” ”Coach,” ”Teacher,” ”Nurse,” ”Therapist,” and ”Librarian,” offering a broader range of design applications. Notably, these Facilitator metaphors tend to be introduced by designers or researchers rather than being directly adopted by users, in contrast to the more user-driven application of Entertainer metaphors. This discrepancy may stem from the fact that Facilitator metaphors are primarily used in specialized domains like health or finance, where authority or structure is crucial. However, these metaphors also raise concerns about assigning social roles to VUIs, which can present ethical and practical challenges (Simpson and Crone, 2022). An alternative approach could involve using non-human or fictional metaphors rather than conventional social roles within the Facilitator category, although finding non-human or fictional metaphors that convey both high formality and high dominance remains a challenge.
A potential solution to these challenges is the use of a layered metaphor strategy. In specialized contexts where accuracy and authority are critical, the social metaphors within the Facilitator category can establish a foundational persona, setting clear expectations for users. Rather than restricting the VUI’s persona to a single metaphor, designers can adopt a more dynamic approach, integrating multiple metaphors that align with specific use-contexts. This is akin to how the desktop metaphor is augmented by concepts such as files, folders, and icons in graphical user interfaces.
This approach offers two advantages. First, it allows for a consistent overarching identity. Second, it introduces adaptability through secondary metaphors that match the nuances of particular interactions—whether that requires the playful supportiveness of Entertainers or the neutral assistance of Guides. For example, in a health-focused VUI, the primary metaphor presented to the user could be that of a ”Nurse,” embodying care and trust. However, context-specific layers of interaction might utilize a ”Search Engine” metaphor to manage conversational recovery after an error or an ”Encyclopedia” metaphor when the user requests in-depth information on a topic. By employing this layered metaphor strategy, designers can create rich, contextually appropriate experiences that align with user expectations while maintaining ease of interaction.
The success of this solution lies in thoughtful execution. Designers must carefully map out intents and contexts within the VUI, assigning appropriate metaphors that enhance the interaction without overshadowing the foundational metaphor. In practice, this involves user studies to determine effective metaphor pairings and iterative design processes to refine how these metaphors are presented and interacted with.
@@ 5. Study 2: Comparing Metaphor-Fluid VUI to a Default VUI
Informed by the results of RQ1, we designed a within-subjects study (N=91) comparing a Metaphor-Fluid VUI and a Default VUI to answer RQ2. The Metaphor-Fluid VUI used four distinct metaphors identified as the most appropriate for use-contexts: Commands, Information Seeking, Sociality, and Error Recovery. Specifically, Genie metaphor was used for Commands, “Computer” from Star Trek for Information Seeking, Admirer for Sociality, and Search Engine metaphor for Error Recovery. In contrast, the Default VUI maintained a single, consistent Assistant metaphor across all interactions and was designed to emulate Google Assistant. This allowed us to examine the effects of metaphor-fluidity on perceived enjoyment, intention to adopt, trust, likability, and intelligence in VUI interactions.
@@ 5.1.1. Designing Metaphorical VUIs
In our study, we designed two VUIs: a Metaphor-Fluid VUI and a Default VUI. These designs fall within the broader domain of conversation design, a field that is still in its formative stages. Currently, there is a notable lack of standardized design processes followed by conversation designers (Sadek et al., 2023). Moreover, there are few design tools or pre-defined personas that can be readily deployed, leaving designers to navigate these challenges on their own. As a result, conversation designers often rely on guidelines provided by industry leaders such as Google and IBM, adapting these frameworks to suit specific contexts. To address these challenges, the first author, who has eight years of experience in conversation design, designed the interactions for our study. Two experts, each with six years of experience, further validated the designs, ensuring an informed approach to the development of the VUIs.
While designing the VUI interactions, our primary focus was to ensure ecological validity. To achieve this, we drew on insights from Bentley et al. (Bentley et al., 2018), who analyzed user logs from 88 diverse homes over a 110-day period, encompassing 66,459 interactions. Their study explored the long-term use of voice assistants, revealing two key patterns that informed our design: (1) users fluidly transitioned between use-contexts within the same interaction session, and (2) the average session length was 5.4 interaction turns. Incorporating these findings, we developed scripts featuring six interaction turns, including five unique commands and one error recovery instance, across four use-contexts.
@@ 5.1.2. Designing for Voice
The audio clips for this study were generated using the text-to-speech (TTS) software, Speechify. [^4] https://speechify.com/ Speechify was chosen for its ability to produce realistic synthetic conversations using multiple voices, which aligned with the study’s design requirements. While an initial approach considered employing a human voice actor for the user’s role and reserving TTS exclusively for the VUI, pilot testing revealed challenges in maintaining a consistent pitch and tone with the human voice across the experimental conditions. These inconsistencies risked introducing confounding variables, as prior research, such as Dubiel et al. (Dubiel et al., 2024b, 2020), has shown that subtle differences in vocal characteristics can significantly affect user perceptions. Consequently, we opted to use synthetic voices for both the user and the VUI.
In choosing the synthetic voices, we aimed to align with gender representation norms observed in commercial VUIs, such as Siri and Alexa, which often use female-sounding voices (Curry et al., 2020). To ensure a clear auditory distinction between the VUI and the user, we assigned a female-sounding voice to the VUI and a male-sounding voice to the user. This decision was informed by literature suggesting that voice gender may influence user perception (Kuzminykh et al., 2020). To further mitigate biases associated with the VUI’s identity, we designated the VUI as ‘Z,’ a neutral name devoid of gender or racial connotations, to minimize confounding effects on the study’s outcomes. For consistency across all experimental conditions, we used Speechify’s synthetic male voice ”Guy” for the user, setting the tone to ”chat” to enhance the voice’s human-like qualities. For the VUI, we selected the female voice ”Aria” with a neutral tone setting of ”none.” All other parameters, such as speech rate, speed, and volume, were standardized at zero percent modifications to ensure uniformity.
Before each conversation, participants heard a brief introductory prompt: ”You are about to hear a conversation between a user with a male-sounding voice and a voice interface named ‘Z’, with a female-sounding voice.” This prompt was delivered using the synthetic voice ”Davis,” configured with the same settings as the study clips to ensure consistency and avoid introducing any additional variables. Additionally, while metaphors were embedded in the design of the VUI, their presence was not made explicit to participants. This decision was informed by prior literature, which found no significant effect of user awareness of metaphors on their perceptions of VUIs (Desai et al., 2024a).

### v2

@@ 3.1.1. Metaphor Identification
Our metaphor collection drew from two sources combining literature review and user studies. Desai & Twidale (Desai and Twidale, 2023) conducted a systematic literature review of VUI research from the past decade, screening 664 articles down to 80, plus a user study with 14 participants interacting with Amazon Echo Dot through semi-structured interviews. Together, these identified 93 VUI-related metaphors (Source A in Figure 2). Chin et al. (Chin et al., 2024) contributed through a study with 58 participants in a simulated smart home environment, where participants completed cognitive tasks and were explicitly asked to compare the VUI to other entities using metaphors, yielding 225 metaphors (Source B in Figure 2).
Figure 2. A total of 93 metaphors were identified from Source A ((Desai and Twidale, 2023)), and 225 metaphors were collected from Source B ((Chin et al., 2024)). After removing duplicates, 98 unique metaphors remained. From these, 20 metaphors were selected based on their frequency of occurrence and feasibility for practical implementation in VUI design.
@@ 3.1.2. VUI Conversational Use-Contexts
Four conversational use-contexts were previously identified in prior literature (Desai and Twidale, 2023)—commands, information seeking, sociality, and error recovery—based on different types of tasks users perform. Although error recovery is not a task itself, it leads to conversations and evokes metaphorical comparisons from users (e.g., silly child, pet). We adopt Epley et al. (2007)’s definition of sociality to include social tasks such as games and small talk. While Desai and Twidale (2023) classified use-contexts through metaphor analysis, other studies used usage logs to categorize conversational use-contexts. Bentley et al. (2018) analyzed logs from 88 households over 110 days, identifying ten use-contexts: music, information, automation, small talk, alarms, weather, video, time, lists, and others. Kim and Choudhury (2021) conducted a 16-week study with 12 older adults, identifying eight topics: music, search, basic device control, casual conversation, time, reminders, weather, and others. Ammari et al. (2019) analyzed 82 Alexa and 88 Google Home devices, identifying eight themes: search, music, timers, automation, smart home, macros, family interaction, and privacy. These findings align with the metaphorical use-contexts (Desai and Twidale, 2023). Commands encompass task types like music, automation, alarms, weather, timers—sharing a one-shot call-and-response pattern where users expect immediate responses. Information seeking includes search-related tasks across studies (Ammari et al., 2019; Bentley et al., 2018; Kim and Choudhury, 2021), while sociality encompasses small talk and casual conversation. Error recovery represents a distinct use-context because errors persist in commercial VUIs—Kim and Choudhury (2021) reported 33% error rates—requiring back-and-forth clarification following unique conversational rules (Porcheron et al., 2018; Fischer et al., 2019)
@@ 3.2. Measures
To determine the most suitable metaphor aligned with user preferences, we developed a methodology that maps metaphors on a 7-point Likert scale across dimensions of hierarchy and formality. We then assessed the desired levels of hierarchy and formality for VUIs in previously identified conversational use-contexts. It is crucial to emphasize that our study’s objective is to identify metaphors that best match users’ mental models for specific VUI tasks. While the metaphors we identify may have broad applicability, designers should ultimately select metaphors that align with their specific design goals.
@@ 3.4. Participants
Our study initially recruited 160 participants through Prolific, [^3] https://www.prolific.com/, a research recruitment platform. However, 30 participants were excluded due to failing attention check questions or timing out, leaving a final sample of 130 participants with an average age of 38.19 years (SD=12.79 years). Selection criteria included a 95% approval rate and at least one year of platform activity. We sought individuals experienced with voice interfaces to ensure meaningful engagement with the study’s tasks and scenarios. To control for cultural variations in metaphor interpretation (Moser, 2000), we limited participation to English-proficient individuals based in the U.S. The participant pool comprised 50.8% females, 46.2% males, and 3.1% non-binary individuals, with 70% holding at least an Associate’s degree. While all participants were U.S. residents and fluent in English, 97% were native speakers. Regarding VUI familiarity, all participants reported at least moderate familiarity, with 75% claiming high or very high familiarity. Usage frequency varied, with all participants using VUIs weekly and 23.1% engaging daily. The survey took an average of 8.5 minutes to complete, and participants received compensation at a fair rate, as informed by Prolific’s policies.
@@ 3.5.1. Mapping Metaphors
Our analysis includes data from all five task scenarios to determine the average formality and hierarchy preferences for each context. The results revealed distinct preferences across different interaction types. When seeking information, participants showed a preference for VUIs to exhibit the highest levels of formality (M=4.67) and dominance (M=3.68). In contrast, during social interactions, participants favored a less formal approach from VUIs (M=2.64). Notably, in error recovery situations, participants expressed a preference for VUIs to display the least dominance (M=2.90) while maintaining a slightly formal demeanor (M=4.49). For command-based interactions, participants desired VUIs to maintain neutral formality (M=3.88) and adopt a slightly subordinate behavior (M=3.06). Considering all four use-contexts collectively, the ideal VUI, according to participants, should demonstrate neutral formality (M=3.92) and a slightly subordinate stance (M=3.23). These findings suggest a nuanced preference for VUI behavior that adapts to the specific nature of the interaction. Table 2 summarizes the distribution for formality and hierarchy ratings.
@@ 3.6. Discussion
Out of our total selection of 20 metaphors, 5 metaphors were non-human—such as ”Computer” from Star Trek, Encyclopedia, Family Pet, Genie, and Search Engine—and 2 metaphors, specifically ”Computer” from Star Trek and Genie, were fictional in nature. Out of all use-contexts, interestingly, only one human metaphor—Admirer for Sociality—was rated as the most suitable for the chosen use-contexts.
This finding is insightful because it challenges the common assumption that human metaphors would naturally be the most relatable or suitable for VUIs. Instead, our results suggest that non-human and fictional metaphors are often more compatible with users’ desired levels of formality and hierarchy in interactions. This aligns with previous work, such as Pradhan et al. (Pradhan et al., 2019), who found that older adults categorize VUIs as both persons and objects, underscoring the unique hybrid role VUIs occupy. Similarly, Jung et al. (Jung et al., 2022) found no significant preference for human over non-human metaphors in chatbots, challenging the assumption that human-like qualities are inherently more effective. Desai et al. (Desai et al., 2024b) further showed that metaphor preferences are context-dependent, with human metaphors preferred in health but no clear preference in finance.
These findings provide important implications for VUI designers: selecting metaphors that users find appropriate may require moving beyond the human metaphor and instead embracing characters that exist outside of human conventions and cultural associations. There is a natural tendency in conversation design to rely on human metaphors (Desai and Twidale, 2023; Doyle et al., 2019), though these often come with ethical and practical concerns, as previously discussed. By contrast, non-human metaphors remain relatively underutilized, even though previous findings suggest that these metaphors may be better suited for user-centered interaction design (Desai et al., 2024b).
@@ 4. Study 2: Comparing Metaphor-Fluid VUI to a Default VUI
Informed by RQ1’s findings, Study 2 (N=91) compared a Metaphor-Fluid VUI against a Default VUI using a within-subjects design. The Metaphor-Fluid VUI employed the optimal metaphors identified for each use-context: Genie for Commands, ”Computer” from Star Trek for Information Seeking, Admirer for Sociality, and Search Engine for Error Recovery. The Default VUI maintained Google Assistant’s consistent Assistant metaphor across all interactions. This design enabled examination of how metaphor-fluidity affects perceived enjoyment, intention to adopt, trust, likability, and intelligence in VUI interactions.
@@ 4.1.1. Designing Metaphorical VUIs
In our study, we designed two VUIs: a Metaphor-Fluid VUI and a Default VUI. These designs fall within the broader domain of conversation design (Sadek et al., 2023). The VUIs were designed by the first author, who has eight years of experience in conversation design, and validated by two experts, each with six years of experience, ensuring an informed approach to the development of the VUIs.
@@ 4.1.2. Designing for Voice
The audio clips for this study were generated using the text-to-speech (TTS) software, Speechify. [^4] https://speechify.com/ Speechify was chosen for its ability to produce realistic synthetic conversations using multiple voices, which aligned with the study’s design requirements. While an initial approach considered employing a human voice actor for the user’s role and reserving TTS exclusively for the VUI, pilot testing revealed challenges in maintaining a consistent pitch and tone with the human voice across the experimental conditions. These inconsistencies risked introducing confounding variables, as prior research, such as Dubiel et al. (Dubiel et al., 2024, 2020), has shown that subtle differences in vocal characteristics can significantly affect user perceptions. Consequently, we opted to use synthetic voices for both the user and the VUI.
In choosing the synthetic voices, we aimed to align with gender representation norms observed in commercial VUIs, such as Siri and Alexa, which often use female-sounding voices (Curry et al., 2020). To ensure a clear auditory distinction between the VUI and the user, we assigned a female-sounding voice to the VUI and a male-sounding voice to the user. This decision was informed by literature suggesting that voice gender may influence user perception (Kuzminykh et al., 2020). To further mitigate biases associated with the VUI’s identity, we named the VUI ‘Z,’ a neutral name devoid of gender or racial connotations, to minimize confounding effects on the study’s outcomes (Desai et al., 2024a). For consistency across all experimental conditions, we used Speechify’s synthetic male voice ”Guy” for the user, setting the tone to ”chat” to enhance the voice’s human-like qualities. For the VUI, we selected the female voice ”Aria” with a neutral tone setting of ”none.” All other parameters, such as speech rate, speed, and volume, were standardized at zero percent modifications to ensure uniformity.
Before each conversation, participants heard a brief introductory prompt: ”You are about to hear a conversation between a user with a male-sounding voice and a voice interface named ‘Z’, with a female-sounding voice.” This prompt was delivered using the synthetic voice ”Davis,” configured with the same settings as the study clips to ensure consistency and avoid introducing any additional variables. Additionally, while metaphors were embedded in the design of the VUI, their presence was not made explicit to participants. This decision was informed by prior literature, which found no significant effect of user awareness of metaphors on their perceptions of VUIs (Desai et al., 2024b).